OpenStudy (anonymous):

PLZ HELP solve for p 3 (p+q) = p help me

OpenStudy (anonymous):

you have to distribute the three to the p and q

OpenStudy (anonymous):

So like 3p + 3q = p?

OpenStudy (anonymous):

exactly then you take 3p to the other side and divide to get 1/3p

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

so divide 3p to p?

OpenStudy (anonymous):

yes that is how you get the one third

OpenStudy (anonymous):

so divide each side by 3p or just p

OpenStudy (anonymous):

\[3q =\frac{ 1 }{ 3 }p\]

OpenStudy (anonymous):

is what you are left with

OpenStudy (trantom):

@kitkatcandy 3(p+q)=p First use the distributive property. 3(p+q) becomes 3p+3q 3p + 3q = p Now subtract p from both sides 2p + 3q = 0 Now subtract 3q from both sides. Instead of saying 0-3q, we can just say -3q 2p = -3q Now divide both sides by 2 p = -3/2q

OpenStudy (anonymous):

Thank you so much.